OpenStudy (anonymous):

When you square each side of an equation, is the resulting equation equivalent to the original?

OpenStudy (anonymous):

not quite this process results in adding one more solution for example, x=2 (2 is the solution) now square : x^2 = 4 (-2,2 are the solutions)

OpenStudy (anonymous):

(adding one more solution in the general case ofcurse) if we had x=0 the square thing wont change anything

OpenStudy (anonymous):

so is the answer yes?

OpenStudy (anonymous):

since it doesnt change?

OpenStudy (anonymous):

but i just showed you that it changes!

OpenStudy (anonymous):

to x=2 we get x^2=4 and we got another solution which is not the solution to the original one

OpenStudy (anonymous):

"the square thing wont change anything" lol confusing me!

OpenStudy (anonymous):

it was just for the case of x=0 read my first comment

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

need more help with something else ?

OpenStudy (anonymous):

most likely yes lol just give me a moment